This is a wave problem
an fm radio station broadcasts at a frequency of 90 mhz. what distance does a radio wave leaving the antenna travel in one vibration if the wave travels at a speed of 3.0 x 10 to the 8th m/s?
1
Expert's answer
2012-11-12T10:43:20-0500
In one vibration, distance isexactly one wavelength, by definition.
Wavelength = speed / frequency.

If you meant mHz, millihertz,
wavelength = (3.0*10^8 m/s) / (90*10^-3 Hz) = 3.33 * 10^9 meters

But by common sense, it's MHz, megahertz, so
wavelength = (3.0*10^8 m/s) / (90*10^6 Hz) = 3.33 meters

Answer: 3.33 meters